# Learned Contextual LiDAR Informed Visual Search in Unseen Environments

### v1

@@ IV-C Hardware Experiments
Experiments are in a 20mx30m indoor apartment setting. The environment is out of distribution as it is not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. The platform used is a Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera. The search target of interest is a small rolling suitcase and YOLOv5 [16] is used for detection of the object. On detection, the exploration node is shut down and the task is complete. We perform several trials over varying initial robot position and target locations. A time lapse of one experiment is shown in Fig. 6 where the search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the one on the left hand side. However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments can be found on the project website https://sites.google.com/view/lives-icra-2024/home.
@@ V Conclusions
We present a method for visual search in unseen environments using contextual LiDAR information to inform planning. Results show LIVES to outperform the closest of several baseline methods in varied environments by 10-30%. Ablative studies on the map-free scan classifier are performed to validate the choice in architecture and training scheme. Furthermore, the proposed approach is deployed on a real robot in an unseen environment. Experiments certify both the visual search planner and the learning method for map-free scan classification in real-world scenarios. Ongoing work aims to extend the LIVES planner to multiple robots.

### v3

@@ IV-C Hardware Experiments
Experiments take place in a 20mx30m indoor apartment setting and a 15mx38m hallway setting, shown in Fig. 5. The environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. A Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera is utilized. The search target of interest is one or two small rolling suitcases and YOLOv5 [18] is used for detection of the object. Objects are localized by first converting PointCloud data into the optical frame and then by conversion to pixel coordinates using camera intrinsics. A time lapse of one experiment in the apartment is shown in Fig. 6 where the search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the one on the left hand side. However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby. Finally, we compare the performance of the trained classifier in these out of distribution environments to ground truth information. The scan classifier accuracy is also evaluated in the two out of distribution experimental environments. The results show $84.0168\%$ accuracy in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment. These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home.
@@ V Conclusions
We present a method for visual search in unseen environments using contextual LiDAR information to inform planning. Results show LIVES to outperform the closest of several baseline methods in varied environments by 10-30%. The proposed approach is deployed on a real robot in an unseen environment to certify the learning method and visual planner for real-world scenarios. Ongoing work aims to extend the LIVES planner to multiple robots.